OpenStudy (anonymous):

Betsey, a recent retiree, requires $6,000 per year in extra income. She has $50,000 to invest and can invest in b-rated bonds paying 15% per year or in a certificate of deposit (cd) paying 7% per year. How much money should be invested in each to realize exactly $6,000 in interest per year?

OpenStudy (amistre64):

amount@15% = (6000 - 50000(.07))/(.15 - .07); i think

OpenStudy (amistre64):

31250(.15) + (the rest)(.07) = 6000 ??

OpenStudy (anonymous):

Which part is correct? I got a different answer?

OpenStudy (amistre64):

31250 @ 15% is correct; which means 18750 @ 7% is left

OpenStudy (amistre64):

you need 2 amounts; a and b a+b total to 50,000; a+b = 50000 a @ 7% + b@15% = 6000, to wit ................................................. a +b = 50000 ; *-.07 a(.07) + b(.15) = 6000 -a(.07) -b(.07) = -50000(.07) a(.07) + b(.15) = 6000 ------------------------------ b(.15-.07) = (6000 - 50000(.07)) b = (6000 - 50000(.07))/(.15-.07)